OpenStudy (anonymous):

Write the equation y - 2 = 2(x - 3) in standard form. A. y - 2x = 4 B. 2x - y = -6 C. 2x + y = 6 D. 2x - y = 4

OpenStudy (anonymous):

y - 2 = 2(x - 3) multiply out the 2 on the right side: y-2=2x-6 bring the y over to the right by subtracting to both sides: y-2-y=2x-6-y -2 = 2x-y-6 add 6 to both sides: -2+6=2x-y-6+6 4=2x-y